Find an equation of a line that passes through points (6,5) and is perpendicular to 2y+3x=6
Answer by robertb(5830)

The equation of the unknown line must have the form 3y - 2x = c. Find c.
Then 3*5 - 2*6 = c, or 15 - 12 = c, or c = 3.
Therefore the equation of the line is 3y - 2x = 3.